OpenStudy (ksaimouli):

f(x)=x^3+4 state any restrictions on the domain ?

OpenStudy (anonymous):

Domain is R

OpenStudy (anonymous):

R being all real numbers

OpenStudy (ksaimouli):

by the way how u know the domain by just looking at the equation plz say

OpenStudy (anonymous):

No you don't If you draw the graph x^(3)+4 =y then you will notice that the domain includes all real numbers

OpenStudy (ksaimouli):

ok thnx

OpenStudy (anonymous):

There are a few things you need to know about domains

OpenStudy (anonymous):

Want me to go over it with you?

OpenStudy (anonymous):

If you have a rational function (a fraction) like x^3/x-3, the domain can't include 3 because the denominator can never be 0 (you cant divide by zero). For a square root function such as (3x+3)^(1/2) or \[\sqrt{ (3x+3}\] you cant take the square root of a negative number thus the domain is restricted to numbers that make 3x+3 positive or zero If you have an equation with that doesn't have either of these things the domain will be R

OpenStudy (anonymous):

all real numbers I mean or \[\mathbb{R}\]

OpenStudy (ksaimouli):

ook thnx u

OpenStudy (anonymous):

You might also have to deal with logarithmic functions

OpenStudy (anonymous):

if you want me to explain them I can

OpenStudy (anonymous):

really domains are easy :P once you understand them

OpenStudy (ksaimouli):

sure

OpenStudy (anonymous):

So look at the definition of \[\log _{b}(x) = y\] which means the same thing as \[b^y = x\] if you draw a logarithmic function you can notice that x can never be 0 or negative Thus the domain is restricted to all numbers that are positive but not zero Same thing with ln function which is just \[\log _{e}(x)\] e being just a number you can look it up

OpenStudy (ksaimouli):

ok thnx aloy